Let A = f(r) rˆ. Verify Gauss’s theorem for the sphere r = R.
Answer #1

Given

vec A=f(r)hat r

Therefore

vec abla.vec A=rac{1}{r^2}rac{d}{dr}(r^2f(r))

Hence

JT dr

Since the radius of the sphere is R, the above integral over r is from 0 to R which gives the result

int_Vvec abla.vec A, dV=4pi R^2 f(R) ...(1)

And

egin{align} oint_Svec A.hat r,dS=oint_S f(r)dS onumber end{align}

where the integral over the surface element is given by

oint dS=r^2oint dOmega=4pi r^2

The surface is at r = R. Therefore

oint_Svec A.hat r,dS=R^2f(R)oint dOmega=4pi R^2f(R) ....(2)

Since equation (1) and (2) are the same, Gauss' divergence is proved.